# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### v1

@@ 1. Introduction
Obviously, CAT represents a complex fusion of machine intelligence and assessment techniques. It needs to manage large question banks, adapt to varying examinee proficiencies, and real-time decision-making. Moreover, practical CAT also involves ensuring reliability, fairness, search efficiency, etc. These challenges make CAT a multifaceted decision-making problem that requires sophisticated solutions. Particularly, with the emergence of large-scale and diverse online testing platforms, traditional statistical methods often struggle to optimally select questions while addressing these multifarious factors. Machine learning (Jordan and Mitchell, 2015), with its ability to learn from data and make precise predictions, offers a viable solution to enhance test efficiency and accuracy within this volatile environment. Previous CAT surveys (Chang, 2015; Vie et al., 2017; Cheng, 2008) have largely revolved around the domains of statistics and psychometrics. Given CAT’s interdisciplinary nature, this paper seeks to understand and review methodologies from a machine-learning perspective, which is more approachable to a wider readership and sheds light on building strong CAT systems.
@@ 2. Background of CAT
  • Healthcare. In healthcare, particularly in mental health, there’s a crucial need for quick and accurate assessments. CAT meets this demand by offering quick evaluations akin to those conducted by skilled clinicians, thereby enhancing the overall efficiency of medical care. CAT for mental health assessments creates scales, including those for depression, anxiety, and suicide risk (Gibbons et al., 2018; Gibbons and deGruy, 2019; Gibbons et al., 2017, 2013, 2014), have shown promising results compared to traditional methods (Kroenke et al., 2001; Bagby et al., 2004).
@@ 3.1. Task Formalization
As illustrated in Figure 2, at test step $t\in[1,2,...,T]$ in CAT, examinee’s current proficiency estimate $\hat{\theta}^{t}$ is estimated using previous $t$ responses; then leverage $\hat{\theta}^{t}$ to retrieve the next question $q_{t+1}$ from question bank $\mathcal{Q}$ to ask examinee, and receive his/her next response $y_{t+1}$. These interactions form a response sequence $\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{T},y_{T})\}$, where $y_{t}=1$ if the response to $q_{t}$ is correct and 0 otherwise. To achieve the goals of CAT, the following two steps are necessary at each test step:
Step 1: Proficiency Estimation. The Cognitive Diagnosis Model (CDM), denoted by $f(\cdot)$, acts as a user model, predicting the probability of a correct response by an examinee with proficiency $\theta$, which is denoted as $f(q,\theta)=P(y=1|q,\theta)$. The implementation of CDM often draws upon cognitive science (De La Torre, 2009) or psychometrics(Ackerman et al., 2003). To accurately estimate his/her proficiency at each step, various estimation methods can be used, e.g., Maximum Likelihood Estimation (MLE) or Bayesian Estimation. In applications, the binary cross-entropy loss is frequently utilized: at step $t$, given previous $t$ responses $X_{1:t}=\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{t},y_{t})\}$, the corresponding empirical loss is:
@@
thus the current estimate of proficiency, $\hat{\theta}^{t}$, is obtained by minimizing the loss function $L(\theta)$: $\hat{\theta}^{t}=\mathop{\arg\min}_{\theta}{{L}(\theta)}$.
Step 2: Question Selection. The heart of CAT is an algorithm that picks the next question $q_{t+1}$ from the question bank $\mathcal{Q}$, using examinee’s current proficiency estimate $\hat{\theta}^{t}$ as a guide:
@@
where $\mathcal{V}_{q}(\hat{\theta}^{t})$ is the value of question $q$. For instance, $\mathcal{V}$ might be a measure of how much information the question will provide about the examinee’s proficiency, or it could be the output of a policy $\pi$ specifically designed to determine question selection.
After receiving new response $y_{t+1}$, CDM updates and estimates proficiency $\hat{\theta}^{t+1}$. The above process will be repeated for $T$ times, ensuring the final step estimate $\hat{\theta}^{T}$ close to the true $\theta_{0}$, i.e.,
@@
However, solving this optimization problem directly is impractical, as the true proficiency $\theta_{0}$ is not observable, and even the examinees may not know their exact proficiency level. Consequently, existing methods are all approximations of this target. For instance, traditional statistical selection methods (Lord, 2012; Chang and Ying, 1996) utilize the asymptotic statistical properties of the MLE to reduce estimation uncertainty, such as choosing questions that closely match the examinee’s current estimated proficiency $\hat{\theta}^{t}$. More recent Subset Selection approaches (Zhuang et al., 2023) try to identify a theoretical approximation of $\theta_{0}$ to serve as a new objective for optimization. For further details, refer to Section 5. Meanwhile, as a practical system, considerations extend beyond this proficiency estimation objective (Definition 1). Factors such as CAT’s question exposure control, robustness, fairness, and search efficiency must be addressed as well. An exhaustive discussion of these factors is presented in Section 7.
Evaluation Methods: To validate the accuracy of the estimated proficiency, two primary approaches are employed: 1) using the examinee’s estimated values within the CDM to predict the correctness label $y$ of the responses on his/her reserved response data, often measured by cross-entropy; 2) using simulation to generate true proficiency values $\theta_{0}$, simulating the examinee’s responses to each question. Then the Mean Squared Error (MSE) between the estimates and the simulated true values can be calculated. The details can be found in Section 8.
@@
 NLP (Yin et al., 2019; Ning et al., 2023). , leaf, text width=22em ] ] [ Question Bank Development (6.2), fill=yellow!10, text width=20em [  Expert-based (Reckase, 2010; He and Reckase, 2014; Van Der Linden et al., 2000; Way et al., 2005; van der Linden et al., 2006; Ariel et al., 2004; Gulliksen, 2013; Swanson and Stocking, 1993)., leaf, text width=22em ] ] ] [ Test Control 7
@@ 4.2. Diagnostic Classification Model
where $s_{j}$ is the slip parameter, indicating the likelihood of an incorrect response despite mastery, and $g_{j}$ is the guess parameter, reflecting the chance of a correct guess in the absence of mastery. Its extension G-DINA(de la Torre, 2011) provides a granular view of examinee proficiency, while FuzzyCDF(Liu et al., 2018) leverages fuzzy set theory for nuanced diagnostics from both objective and subjective data. Another approach, the Attribute Hierarchy Method (AHM)(Leighton et al., 2004), applies rule space theory to structure knowledge dependencies and align examinee proficiencies with the nearest ideal cognitive patterns to obtain diagnostic results.
Compared to Latent Trait Models, Diagnostic Classification Models offer a more granular and comprehensive assessment of examinee proficiencies. They are particularly adept at providing detailed feedback on an individual’s strengths and weaknesses across multiple knowledge concepts, which is important for CAT and targeted further interventions. These models underscore a critical shift towards a more nuanced understanding of learning and proficiency, recognizing the multifaceted nature of knowledge acquisition and adaptive testing.
@@ 5.1.1. General Framework
where $\mathcal{I}_{j}(\cdot)$ is the informativeness of question $q_{j}$ (e.g., Fisher information). As illustrate in Definition 1, CAT assumes that each examinee has a true proficiency value ($\theta_{0}$) and it is considered as a parameter estimation process. The informativeness of a question can thus be interpreted as the expected contribution of the response on this question to the parameter estimation. This concept will be reflected in various selection algorithms discussed later.
Fisher Information. In the parameter estimation problems, Fisher Information (Rissanen, 1996) is a concept from information theory and statistics that measures the amount of information that an observable random variable carries about an unknown parameter upon which the probability of that random variable depends. In the context of CAT, Fisher Information is often used to quantify the amount of information that a question provides about an examinee’s proficiency (Cheng, 2008). Specifically, we consider a random variable $X_{j}=(q_{j},y_{j})$ for which the pdf or pmf is $f(q_{j},\theta)$, where $\theta$ is the unknown parameter. The fisher info contained in the random variable $X_{j}$ is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}[(\nabla_{\theta}L(X_{j}|\theta))^{2}]=\frac{(\nabla_{\theta}f(q_{j},\theta))^{2}}{f(q_{j},\theta)(1-f(q_{j},\theta))}$, where $L(X|\theta)={y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))}$ is the likelihood function of $X$ with respect to the parameter $\theta$.
Thus, when using 3PL-IRT to model the $f(q,\theta)$ and given current estimate $\hat{\theta}^{t}$, the Fisher Information of question $j$ can be calculated as:
@@
Obviously, it can be obtained that as the number of test questions ($t$) or the Fisher information ($\mathcal{I}$) increases, the variance decreases. Since the estimated value is asymptotically unbiased (i.e., its mean equals the true value $\theta_{0}$), a decrease in its variance results in the distribution becoming more “tight” (higher $p(\hat{\theta}^{t}=\theta_{0})$), thereby reducing the uncertainty of the estimated proficiency $\hat{\theta}^{t}$. Due to the unknown $\theta_{0}$, maximizing the Fisher Information (MFI) at the current estimate $\hat{\theta}^{t}$ essentially serves to lower the variance of the proficiency estimate and reduce uncertainty, thereby improving the estimation efficiency. MFI has been widely popular in the development of personalized testing over the decades and extensively applied in practice as well. However, the effectiveness of MFI is most pronounced when the estimated value is close to the unknown true value $\theta_{0}$ (Chang, 2015). Therefore, using the current estimate in place of $\theta_{0}$ can sometimes lead to inaccuracies if the current estimate is far from the true value.
@@
Fisher Information is a mathematical concept that exists independently of CAT. Notably, this concept finds a natural extension in Multidimensional IRT (MIRT), where it transitions from a real number to a matrix (Hooker et al., 2009), facilitating the assessment of multiple abilities or traits simultaneously. Specifically, the information matrix provided by question $j$ at proficiency $\theta$ (now a vector) is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}\left[\nabla\log L(X_{j}|\theta)\nabla\log L(X_{j}|\theta)^{T}\right]$. This Fisher Information matrix is instrumental in MIRT-based personalized assessment systems, as its inverse corresponds to the covariance matrix for maximum likelihood proficiency estimates. Researchers have exploited this relationship to develop many question selection strategies such as D-Optimization, A-Optimization, and E-Optimization, which respectively focus on maximizing the matrix’s determinant, minimizing the trace of its inverse, and maximizing its minimum eigenvalue (Reckase, 2006). These strategies, complemented by Bayesian optimization variants, have considerably advanced the statistical techniques available for question selection in adaptive testing.
Kullback-Leibler Information. Fisher information is a commonly used measure in CAT due to its theoretical foundation, simplicity in mathematical form, and computational efficiency. However, the application of Fisher information might not be as efficient as expected if the estimators are not close to the true value $\theta_{0}$ (Chang, 2015). This issue becomes particularly evident during test’s initial stage, when the test length (number of questions) is too short to provide an accurate proficiency estimate. Consequently, relying on the MSI may not yield a satisfactory results, because Fisher information, often referred to as local information, only considers the information at a specific point ($\hat{\theta}^{t}$).
To address this issue, Chang and Ying (1996) proposed a global information measure based on Kullback-Leibler (KL) divergence, also known as KL information: Let $\theta_{0}$ be the true proficiency, for any $\theta$, the KL information of the question $q_{j}$ (with response $X_{j}$) which measures the discrepancy between the two probability distributions specified by $\theta_{0}$ and $\theta$, is defined by
@@ 5.1.2. Advances and Generalization
In addition to the two aforementioned kinds of information which are representative and popular, we briefly introduce recent statistical works that develop advanced and generalized selection algorithms from an information-based perspective.
@@
Selection Algorithms in CD-CAT. It’s important to note that the selection algorithms mentioned above are all designed based on IRT and are not applicable to other CDMs. From a broader perspective, as discussed in Section 4, there is a wide range of CDMs (e.g., the DINA model (De La Torre, 2009)) that differ from IRT in their approach to characterizing examinee’s proficiency, $\theta$. They represent an examinee’s $\theta$ as a discrete set of knowledge mastery, i.e., cognitive state. The selection algorithms for these specific CDMs are distinguished as Cognitive Diagnosis CAT (CD-CAT) (Cheng, 2009). In CD-CAT, the general idea of designing selection algorithms remains the same: maximizing specific information associated with the selected questions. Techniques grounded in information theory, such as Kullback-Leibler (KL) divergence and Shannon entropy, are pivotal in this pursuit. The KL algorithm (Henson and Douglas, 2005) seeks to enhance discrimination between different latent classes, while Shannon entropy (Tatsuoka, 2002) aims to reduce the uncertainty in estimating an examinee’s latent class.
While continuous traits (as modeled in IRT) and discrete latent states (as modeled in DINA) represent different facets of an examinee’s characteristics, they are not mutually exclusive but rather closely intertwined. This connection has spurred interest in how to assess both simultaneously, leading to the development of dual-objective CD-CAT (Kang et al., 2017; Zheng et al., 2018). This includes methods such as the dual information index (Cheng and Chang, 2007), aggregate ranked information index (Wang et al., 2014), aggregate standardized information (ASI) index (Wang et al., 2014), Jensen-Shannon divergence (Kang et al., 2017), G-DINA discrimination index (Kaplan et al., 2015), dapperness with information index (Dai et al., 2016), and the information product approach (Zheng et al., 2018). Despite the variety of these strategies, their essence remains rooted in the conceptual understanding and quantification of specific statistical information.
@@ 5.2. Active Learning Algorithms
An ideal CAT framework should be model-agnostic, meaning that the selection algorithms can adapt to different CDMs while remaining effective. Researchers are exploring a broader approach to the CAT problem by investigating its correlation with a versatile machine learning technique—Active Learning (Krishnakumar, 2007). Active Learning is to choose valuable data actively, thus can train better models with less data. This method has improved data efficiency and cut annotation costs in numerous learning tasks (Budd et al., 2021).
As shown in Figure 5, in active learning, each cycle begins with a learning model and a strategy for selecting data. The system selects an unlabeled data (set) to label at each iteration, using this to supplement the limited labeled data and improve the model’s performance. The key point behind this is how to design the data selection strategy. Existing methods typically view selection from two perspectives: informativeness and representativeness. Informativeness-based algorithms choose samples that significantly reduce a learning model’s uncertainty (Yoo and Kweon, 2019). Representativeness-based algorithms, on the other hand, choose samples that best reflect the overall patterns of unlabeled data (Ghorbani et al., 2022; Wang et al., 2020b), and their combinations (Ash et al., 2019; Wang et al., 2020b).

### v2

@@ 1. Introduction
Obviously, CAT represents a complex fusion of machine intelligence and assessment techniques. It needs to manage large question banks, adapt to varying examinee proficiencies, and real-time decision-making. Moreover, practical CAT also involves ensuring reliability, fairness, search efficiency, etc. These challenges make CAT a multifaceted decision-making problem that requires sophisticated solutions. Particularly, with the emergence of large-scale and diverse online testing platforms, traditional statistical methods often struggle to optimally select questions while addressing these multifarious factors. Machine learning (Jordan and Mitchell, 2015), with its ability to learn from data and make precise predictions, offers a viable solution to enhance test efficiency and accuracy within this volatile environment. Previous CAT surveys (Chang, 2015; Vie et al., 2017; Cheng, 2008; Mujtaba and Mahapatra, 2020) have largely revolved around the domains of statistics and psychometrics. Given CAT’s interdisciplinary nature, this paper seeks to understand and review methodologies from a machine-learning perspective, which is more approachable to a wider readership and sheds light on building strong CAT systems.
@@ 2. Background of CAT
  • Healthcare. In healthcare, particularly in mental health, there’s a crucial need for quick and accurate assessments. CAT meets this demand by offering quick evaluations akin to those conducted by skilled clinicians, thereby enhancing the overall efficiency of medical care. CAT for mental health assessments creates scales, including those for depression, anxiety, and suicide risk (Gibbons et al., 2018; Gibbons and deGruy, 2019; Gibbons et al., 2017, 2014), have shown promising results compared to traditional methods (Bagby et al., 2004).
@@ 3.1. Task Formalization
As illustrated in Figure 2, at test step $t\in[1,2,...,T]$ in CAT, examinee’s current proficiency estimate $\hat{\theta}^{t}$ is estimated using previous $t$ responses; then leverage $\hat{\theta}^{t}$ to retrieve the next question $q_{t+1}$ from question bank $\mathcal{Q}$ to ask examinee, and receive his/her next response $y_{t+1}$. These interactions form a response sequence $\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{T},y_{T})\}$, where $y_{t}=1$ if the response to $q_{t}$ is correct and 0 otherwise. To achieve the goals of CAT, the following two steps are necessary at each test step: Step 1: Proficiency Estimation. The Cognitive Diagnosis Model (CDM), denoted by $f(\cdot)$, acts as a user model, predicting the probability of a correct response by an examinee with proficiency $\theta$, which is denoted as $f(q,\theta)=P(y=1|q,\theta)$. The implementation of CDM often draws upon cognitive science (De La Torre, 2009) or psychometrics(Ackerman et al., 2003). To accurately estimate his/her proficiency at each step, various estimation methods can be used, e.g., Maximum Likelihood Estimation (MLE) or Bayesian Estimation. In applications, the binary cross-entropy loss is frequently utilized: at step $t$, given previous $t$ responses $X_{1:t}=\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{t},y_{t})\}$, the corresponding empirical loss is:
@@
thus the current estimate of proficiency, $\hat{\theta}^{t}$, is obtained by minimizing the loss function $L(\theta)$: $\hat{\theta}^{t}=\mathop{\arg\min}_{\theta}{{L}(\theta)}$. Step 2: Question Selection. The heart of CAT is an algorithm that picks the next question $q_{t+1}$ from the question bank $\mathcal{Q}$, using examinee’s current proficiency estimate $\hat{\theta}^{t}$ as a guide:
@@
where $\mathcal{V}_{q}(\hat{\theta}^{t})$ is the value of question $q$. For instance, $\mathcal{V}$ might be a measure of how much information the question will provide about the examinee’s proficiency, or it could be the output of a policy $\pi$ specifically designed to determine question selection. After receiving new response $y_{t+1}$, CDM updates and estimates proficiency $\hat{\theta}^{t+1}$. The above process will be repeated for $T$ times, ensuring the final step estimate $\hat{\theta}^{T}$ close to the true $\theta_{0}$, i.e.,
@@
However, solving this optimization problem directly is impractical, as the true proficiency $\theta_{0}$ is not observable, and even the examinees may not know their exact proficiency level. Consequently, existing methods are all approximations of this target. For instance, traditional statistical selection methods (Lord, 2012; Chang and Ying, 1996) utilize the asymptotic statistical properties of the MLE to reduce estimation uncertainty, such as choosing questions that closely match the examinee’s current estimated proficiency $\hat{\theta}^{t}$. More recent Subset Selection approaches (Zhuang et al., 2023) try to identify a theoretical approximation of $\theta_{0}$ to serve as a new objective for optimization. For further details, refer to Section 5. Meanwhile, as a practical system, considerations extend beyond this proficiency estimation objective (Definition 1). Factors such as CAT’s question exposure control, robustness, fairness, and search efficiency must be addressed as well. An exhaustive discussion of these factors is presented in Section 7. Evaluation Methods: To validate the accuracy of the estimated proficiency, two primary approaches are employed: 1) using the examinee’s estimated values within the CDM to predict the correctness label $y$ of the responses on his/her reserved response data, often measured by cross-entropy; 2) using simulation to generate true proficiency values $\theta_{0}$, simulating the examinee’s responses to each question. Then the Mean Squared Error (MSE) between the estimates and the simulated true values can be calculated. The details can be found in Section 8.
@@
 NLP (Yin et al., 2019; Ning et al., 2023). , leaf, text width=22em ] ] [ Question Bank Development (6.2), fill=yellow!10, text width=20em [  Expert-based (Reckase, 2010; He and Reckase, 2014; Van Der Linden et al., 2000; Way et al., 2005; van der Linden et al., 2006; Ariel et al., 2004; Gulliksen, 2013; Swanson and Stocking, 1993)., leaf, text width=22em ] ] ] [ Test Control
@@ 4.2. Diagnostic Classification Model
where $s_{j}$ is the slip parameter, indicating the likelihood of an incorrect response despite mastery, and $g_{j}$ is the guess parameter, reflecting the chance of a correct guess in the absence of mastery. Its extension G-DINA(de la Torre, 2011) provides a granular view of examinee proficiency, while FuzzyCDF(Liu et al., 2018) leverages fuzzy set theory for nuanced diagnostics from both objective and subjective data. Another approach, the Attribute Hierarchy Method (AHM)(Leighton et al., 2004), applies rule space theory to structure knowledge dependencies and align examinee proficiencies with the nearest ideal cognitive patterns to obtain diagnostic results. Compared to Latent Trait Models, Diagnostic Classification Models offer a more granular and comprehensive assessment of examinee proficiencies. They are particularly adept at providing detailed feedback on an individual’s strengths and weaknesses across multiple knowledge concepts, which is important for CAT and targeted further interventions. These models underscore a critical shift towards a more nuanced understanding of learning and proficiency, recognizing the multifaceted nature of knowledge acquisition and adaptive testing.
@@ 5.1.1. General Framework
where $\mathcal{I}_{j}(\cdot)$ is the informativeness of question $q_{j}$ (e.g., Fisher information). As illustrate in Definition 1, CAT assumes that each examinee has a true proficiency value ($\theta_{0}$) and it is considered as a parameter estimation process. The informativeness of a question can thus be interpreted as the expected contribution of the response on this question to the parameter estimation. This concept will be reflected in various selection algorithms discussed later. Fisher Information. In the parameter estimation problems, Fisher Information (Rissanen, 1996) is a concept from information theory and statistics that measures the amount of information that an observable random variable carries about an unknown parameter upon which the probability of that random variable depends. In the context of CAT, Fisher Information is often used to quantify the amount of information that a question provides about an examinee’s proficiency (Cheng, 2008). Specifically, we consider a random variable $X_{j}=(q_{j},y_{j})$ for which the pdf or pmf is $f(q_{j},\theta)$, where $\theta$ is the unknown parameter. The fisher info contained in the random variable $X_{j}$ is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}[(\nabla_{\theta}L(X_{j}|\theta))^{2}]=\frac{(\nabla_{\theta}f(q_{j},\theta))^{2}}{f(q_{j},\theta)(1-f(q_{j},\theta))}$, where $L(X|\theta)={y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))}$ is the likelihood function of $X$ with respect to the parameter $\theta$. Thus, when using 3PL-IRT to model the $f(q,\theta)$ and given current estimate $\hat{\theta}^{t}$, the Fisher Information of question $j$ can be calculated as:
@@
Obviously, it can be obtained that as the number of test questions ($t$) or the Fisher information ($\mathcal{I}$) increases, the variance decreases. Since the estimated value is asymptotically unbiased (i.e., its mean equals the true value $\theta_{0}$), a decrease in its variance results in the distribution becoming more “tight” (higher $p(\hat{\theta}^{t}=\theta_{0})$), thereby reducing the uncertainty of the estimated proficiency $\hat{\theta}^{t}$. Due to the unknown $\theta_{0}$, maximizing the Fisher Information (MFI) at the current estimate $\hat{\theta}^{t}$ essentially serves to lower the variance of the proficiency estimate and reduce uncertainty, thereby improving the estimation efficiency. MFI has been widely popular in the development of personalized testing over the decades and extensively applied in practice as well. However, the effectiveness of MFI is most pronounced when the estimated value is close to the unknown true value $\theta_{0}$ (Chang, 2015). Therefore, using the current estimate in place of $\theta_{0}$ can sometimes lead to inaccuracies if the current estimate is far from the true value. Fisher Information is a mathematical concept that exists independently of CAT. Notably, this concept finds a natural extension in Multidimensional IRT (MIRT), where it transitions from a real number to a matrix (Hooker et al., 2009), facilitating the assessment of multiple abilities or traits simultaneously. Specifically, the information matrix provided by question $j$ at proficiency $\theta$ (now a vector) is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}\left[\nabla\log L(X_{j}|\theta)\nabla\log L(X_{j}|\theta)^{T}\right]$. This Fisher Information matrix is instrumental in MIRT-based personalized assessment systems, as its inverse corresponds to the covariance matrix for maximum likelihood proficiency estimates. Researchers have exploited this relationship to develop many question selection strategies such as D-Optimization, A-Optimization, and E-Optimization, which respectively focus on maximizing the matrix’s determinant, minimizing the trace of its inverse, and maximizing its minimum eigenvalue (Reckase, 2006). These strategies, complemented by Bayesian optimization variants, have considerably advanced the statistical techniques available for question selection in adaptive testing. Kullback-Leibler Information. Fisher information is a commonly used measure in CAT due to its theoretical foundation, simplicity in mathematical form, and computational efficiency. However, the application of Fisher information might not be as efficient as expected if the estimators are not close to the true value $\theta_{0}$ (Chang, 2015). This issue becomes particularly evident during test’s initial stage, when the test length (number of questions) is too short to provide an accurate proficiency estimate. Consequently, relying on the MSI may not yield a satisfactory results, because Fisher information, often referred to as local information, only considers the information at a specific point ($\hat{\theta}^{t}$). To address this issue, Chang and Ying (1996) proposed a global information measure based on Kullback-Leibler (KL) divergence, also known as KL information: Let $\theta_{0}$ be the true proficiency, for any $\theta$, the KL information of the question $q_{j}$ (with response $X_{j}$) which measures the discrepancy between the two probability distributions specified by $\theta_{0}$ and $\theta$, is defined by
@@ 5.2. Active Learning Algorithms
An ideal CAT framework should be model-agnostic, meaning that the selection algorithms can adapt to different CDMs while remaining effective. Researchers are exploring a broader approach to the CAT problem by investigating its correlation with a versatile machine learning technique—Active Learning (Krishnakumar, 2007). Active Learning is to choose valuable data actively, thus can train better models with less data. This method has improved data efficiency and cut annotation costs in numerous learning tasks (Budd et al., 2021). As shown in Figure 5, in active learning, each cycle begins with a learning model and a strategy for selecting data. The system selects an unlabeled data (set) to label at each iteration, using this to supplement the limited labeled data and improve the model’s performance. The key point behind this is how to design the data selection strategy. Existing methods typically view selection from two perspectives: informativeness and representativeness. Informativeness-based algorithms choose samples that significantly reduce a learning model’s uncertainty (Yoo and Kweon, 2019). Representativeness-based algorithms, on the other hand, choose samples that best reflect the overall patterns of unlabeled data (Ghorbani et al., 2022; Wang et al., 2020b), and their combinations (Wang et al., 2020b).